OpenStudy (anonymous):

I need help fast! I need to know how to find the Domain, Range, X-intercept, Y-intercept,the increasing form, the decreasing form,the Max/min and the end behavior. I just need to know how to find these within exponential equations.

OpenStudy (anonymous):

In your typical exponential function, f(x)=a^x, The domain is of course, any real number.\((-\infty,\infty)\) The range, by subbing the domain in to the function, is \((0,\infty)\) There is no X-intercept in this function (seen from range) There is one Y-intercept. f(0)=a

OpenStudy (anonymous):

The increasing form is when it's \(f(x)=a^x\) The decreasing form is when it's \(f(x)=a^{-x}\) The Max/Min you can get from differentiating. End behaviour? This: